OpenStudy (anonymous):

What is the simplified form of the quantity x over 4 minus y over 3 all over the quantity x over 3 plus y over 4?

OpenStudy (anonymous):

@liana1026 help plz

OpenStudy (anonymous):

@liana1026

OpenStudy (anonymous):

4x-3y / 3x+4y

OpenStudy (anonymous):

can you answer another

OpenStudy (anonymous):

sure as long as u can help me on mine after

OpenStudy (anonymous):

What are the discontinuities of the function f(x) = the quantity x squared minus 16 over the quantity 6x minus 24?

OpenStudy (anonymous):

i have a couple more btw

OpenStudy (anonymous):

6x - 24 = 0 6x = 24 x = 24/6 x = 4 makes the denominator zero, so this causes the discontinuity

OpenStudy (anonymous):

What is the graph of the function f(x) = the quantity of x squared plus 3 x plus 5, all over x plus 2?

OpenStudy (anonymous):

hmm idk that one sorry xc

OpenStudy (anonymous):

What is the graph of the function f(x) = the quantity x squared minus 9x plus 20 over the quantity x minus 4?

OpenStudy (anonymous):

do you know this one

OpenStudy (anonymous):

yes one sec

OpenStudy (anonymous):

x^2 - 15x + 36 x^2-12x-3x+36 x(x-12)-3(x-12) (x-3)(x-12) And now the denominator 5(x-12) Which gives us ................................ x-3 over 5

OpenStudy (anonymous):

well whats the graph look like

OpenStudy (anonymous):

how would i put it into a graph

OpenStudy (anonymous):

hmm not sure i would put this and say i didnt understand how to put in a graph

OpenStudy (anonymous):

a line that crosses the x axis at negative 5 and the y axis at 5. Discontinuity exists at negative 4, 1. a line that crosses the x axis at 5 and the y axis at negative 5. Discontinuity exists at 4, negative 1. a line that crosses the y-axis at 5 and has a slope of negative 1. Discontinuity exists at negative 4, 9. a line that crosses the x axis at negative 5 and the y axis at negative 5. Discontinuity exists at 4, negative 9.

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

are you still here @liana1026

OpenStudy (anonymous):

i have one more after this

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

After the fraction x plus 1 all over 2 minus the fraction x plus 2 all over 3 x have been combined using the least common denominator of 6x, what is the numerator?

OpenStudy (anonymous):